OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1/2x − 2, find f[g(−6)].

OpenStudy (anonymous):

When you see f[g(x)], you're basically taking the result of g(x) and using that as the value of X in f(x). g(-6) = (1/2)(-6)-2 = ?

OpenStudy (anonymous):

So, this is what you do. First you substitute -6 for the x 's in the second equation then you substitute what you get there for the x 's in the first equation.

OpenStudy (anonymous):

g(-6) = (1/2)(-6)-2 = -5 f(g(-6)) = -5 - 4 = -9

OpenStudy (anonymous):

so the answer is -9

OpenStudy (anonymous):

I got -5 for the second equation. Then entered that in the first and got -9 for the first equation.